Calculate the molar hydroxide ion concentration of each of the following solutions given the pH: Ph= 4.55
1
Expert's answer
2018-05-25T07:05:40-0400
pOH = 14 - pH = 14 - 4.55 = 9.45
[OH-] = 10^-pOH
[OH-] = 10^-9.45 = 3.55•10^-10 mol/L